Time Machine
Never again. Thirty items in there, and I'm lacking most of them. One toy and 29 wardrobe items and learnies. I got the toy. Happened before. The chances are not even in these machines. Not gonna bother.
Pixie Dice : roll a 2 or a 12
There is a Foxfire Forest goal where you have to roll a 2 or a 12 - i.e., snake eyes or boxcars.
The chance of rolling a 1 on the first die is 1/6. The chance of rolling a 1 on the second die is also 1/6. So the chance of both happening is 1/36.
Similarly, the chance of rolling a 6 on both dice is also 1/36.
The chance of either of those two rolls happening is 1/36 + 1/36 = 1/18
So about every 18 times of playing Pixie Dice, the player should roll either a 2 or a 12.
The game is available every 16 hours, I probably do it once every 20 hours.
So I should win about every 18 * (20/24) = 15 days.
I have been working on that goal for months.
So, loaded dice. It's fun figuring out these things, anyways :slight_smile:
(Only needed for Foxfire goals, not needed to get to Marada goal 75)
Seven Heaven : Win with 6+ Multiplier (roll 5 doubles)
The chance to roll 1 & 1 is 1/6 * 1/6 = 1/36. Same for the other doubles.
The chance to roll any of the 6 possible doubles is 6 * (1/36) = 1/6
The chance to do that 5 times is 1/6 ** 5 = 1/7776
So, on the average, 7776 times. Less actually, since you have 9 rolls in which to get the 5 doubles. Let's say 7776 * 5 / 9 = 4320 times.
You can try every 3 hours, that's 8 tries/day, so that means an average of 4320/8 = 540 days, assuming you play every 3 hours.
IF you do it EVERY THREE HOURS, that's 1.5 years.
But no one is going to play the game EVERY 3 HOURS - since people have lives and sleep, 3 years is probably a better estimate ......
..... ahem .... IF you roll no 7s, ever ....
Yup, there's that part to the game as well. If you roll a 7, your game is over, even if it was your first roll. To get 5 doubles, you have to NOT roll a 7 at least 5 times - and 7 is the most common roll.
On the average, 1/6 of the time, the sum of the dice will be 7, and 5/6 of the time it will not be 7. So the chance of not getting a 7 in the first 5 rolls is (5/6) ** 5 = .402 - and 3
Keep reading →